# STAR: Constraint LoRA with Dynamic Active Learning for Data-Efficient Fine-Tuning of Large Language Models

### v1

@@ Abstract
Though Large Language Models (LLMs) have demonstrated the powerful capabilities of few-shot learning through prompting methods, supervised training is still necessary for complex reasoning tasks. Because of their extensive parameters and memory consumption, both Parameter-Efficient Fine-Tuning (PEFT) methods and Memory-Efficient Fine-Tuning methods have been proposed for LLMs. Nevertheless, the issue of large annotated data consumption, the aim of Data-Efficient Fine-Tuning, remains unexplored. One obvious way is to combine the PEFT method with active learning. However, the experimental results show that such a combination is not trivial and yields inferior results. Through probe experiments, such observation might be explained by two main reasons: uncertainty gap and poor model calibration. Therefore, in this paper, we propose a novel approach to effectively integrate uncertainty-based active learning and LoRA. Specifically, for the uncertainty gap, we introduce a dynamic uncertainty measurement that combines the uncertainty of the base model and the uncertainty of the full model during the iteration of active learning. For poor model calibration, we incorporate the regularization method during LoRA training to keep the model from being over-confident, and the Monte-Carlo dropout mechanism is employed to enhance the uncertainty estimation. Experimental results show that the proposed approach outperforms existing baseline models on three complex reasoning tasks.111 Our code and results will be available at https://github.com/callanwu/STAR.
@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Wei et al. (2021); OpenAI (2022); Touvron et al. (2023a, b); Zhao et al. (2023) have demonstrated the powerful capabilities of zero/few-shot learning with prompting techniques, including In-Context Learning Dong et al. (2022) and Chain-of-Thought Wei et al. (2022), where no parameter update is required. However, previous work Hendrycks et al. (2020); Yuan et al. (2023); Bai et al. (2023); Isik et al. (2024) has shown that further fine-tuning is still crucial for tasks involving complex reasoning such as arithmetic reasoning Roy and Roth (2016); Cobbe et al. (2021), commonsense reasoning Mihaylov et al. (2018); Clark et al. (2019), etc.
@@
Fine-tuning LLMs requires updating a large number of parameters, which takes a lot of time and consumes considerable memory. Taking LLaMA-7B Touvron et al. (2023a) as an example, fine-tuning it on a dataset of 52k instances takes over 24 hours on a Nvidia V100 with over 28GB GPU memory Bommasani et al. (2021). Therefore, Parameter Efficient Fine-Tuning (PEFT) methods  Houlsby et al. (2019); Lester et al. (2021); Li and Liang (2021); Hu et al. (2021); Ding et al. (2023) and Memory Efficient Fine-Tuning (MEFT) methods Liao et al. (2023) have been proposed. In addition to updating a vast number of parameters and consuming substantial memory, a neglected factor in LLMs fine-tuning is the extensive consumption of annotation data. Moreover, due to the inherent complexity of tasks, the human annotation resources required for fine-tuning LLMs are also significant Ouyang et al. (2022).
Therefore, it is important to develop the Data-Efficient Fine-Tuning (DEFT) method for LLMs. A common practice to improve data efficiency is active learning Cohn et al. (1996); Settles (2009), while it has been shown that the PEFT methods can alleviate the reliance on annotated data to some extent Ding et al. (2023). A straightforward idea for DEFT is to combine the PEFT method with active learning. However, such a combination is not trivial. As shown in Figure 1 (a), simply fine-tuning an LLM with LoRA Hu et al. (2021) under an uncertainty-based active learning framework yields consistently inferior performances compared to passive learning (random selection of data in active learning) on three datasets, while as shown in Figure 1 (b), fine-tuning LLM with full parameter under active learning yields performances better than passive learning.
@@ 2.1 Efficient Fine-tuning Methods
As LLMs continue to expand in size, the computational and financial resources required for fine-tuning these models become increasingly prohibitive. To address this challenge, Efficient Fine-Tuning (EFT) has emerged as an essential area of research Wan et al. (2023). EFT methods can be classified into PEFT and MEFT Wan et al. (2023); Liao et al. (2023). PEFT, in particular, aims to adjust a minimal subset of the model’s parameters, thus conserving computational resources while maintaining or enhancing model performance Hu et al. (2023). We classify PEFT methods into four categories: Prompt Tuning Lester et al. (2021); Liu et al. (2023b) only allow an additional $k$ tunable tokens per downstream task to be prepended to the input text. Prefix Tuning Li and Liang (2021); Liu et al. (2022) keeps language model parameters frozen but optimizes a small continuous task-specific vector that pretends to be key-value pairs. Adapter Houlsby et al. (2019); He et al. (2021) is a new module added between layers of a pre-trained network, which is a bottleneck architecture. Low-Rank Adaptation (LoRA) Aghajanyan et al. (2021); Hu et al. (2021) reduces the parameters and enhances computational efficiency by applying low-rank matrices. Moreover, minimizing memory usage in fine-tuning for improving efficiency has also emerged as a critical topic Liao et al. (2023), with several innovative solutions being proposed. Among these, techniques such as QLoRA Dettmers et al. (2023), QA-LoRA Xu et al. (2023), and LoftQ Li et al. (2023) stand out for their ability to significantly reduce memory requirements without compromising model performance. In this paper, we focus on the application of LoRA.
@@ 2.2 Active Learning with LLMs
Active Learning (AL) has been extensively investigated across a multitude of NLP tasks, encompassing machine translation Miura et al. (2016); Zhao et al. (2020), natural language inference Snijders et al. (2023), named entity recognition Shen et al. (2017) and text classification Ein-Dor et al. (2020); Margatina et al. (2022); Schröder et al. (2023). In the era of LLMs, active learning is primarily employed in the selection of prompts and the annotation of data Zhang et al. (2023b); Liu et al. (2023a). For instance,  Margatina et al. (2023) explores various active learning strategies for selecting the most relevant examples for in-context learning with LLMs.  Diao et al. (2023) introduces an active prompting method that leverages uncertainty metrics to select questions for annotation. In the domain of integrating PEFT with AL, Jukić and Snajder (2023a) explored PEFT methods with different active learning in Pre-trained language models (PLMs) such as BERT Devlin et al. (2019), demonstrating that the integration of PEFT with active learning can offer substantial performance gains. Different from Jukić and Snajder (2023a), we apply decoder-only generative LLMs as the backbones, to our knowledge, we are the first to integrate LLMs combined PEFT with AL within the realm of reasoning tasks.
@@ 4.1 Probe Experiment Design
We mainly focus on two training variants of LLMs. The first one, denoted as PEFT method, is LLM finetuned with LoRA, which is shown to be problematic under the AL framework. The second one, denoted as Few-shot method, is untuned LLM, which is taken as a control group. To enhance the performance of untuned LLM on downstream tasks, we employ In-Context Learning Dong et al. (2022) by adding demonstrations with the input prompt. LLaMA-2 Touvron et al. (2023b) serves as the backbone LLM. Experiments are conducted on BoolQ Clark et al. (2019) because the labels of this dataset only include “true” and “false”, which makes the prediction uncertainty and prediction confidence easy to calculate.
@@ 4.2 Probing with Prediction Confidence
where $p_{true}$ and $p_{false}$ denotes the probabilities of token “true” and “false” respectively.
Then the prediction confidence of PEFT and Few-shot are calculated and the density plot is drawn to make a comparison between these two methods. To mitigate differences in model accuracy, we only consider the confidence of the model for the wrong predictions on the test set of BoolQ. The intuition is that for examples that the model is less likely to predict right, they should be less confident.
@@ 4.3 Probing with Prediction Entropy
Figure 3: (a) Heatmap of correlation between prediction entropy across different iterations. (b) Scatter plot for prediction entropy between base model (Iter0) and model after first iteration (Iter1). (c) Same as (b), except values are taken from Iter5 and Iter6.
@@
The correlation between examples with top 1000 entropy at the beginning is calculated and the heatmap of the correlation is shown in Figure 3 (a). As we can observe in Figure 3 (a), the correlation between the base model (model without PEFT tuning) and models after AL iteration is close to 0, which indicates a clear gap between the base model and PEFT model. This phenomenon is even clear with the scatter plot in Figure 3, where the dots in Figure 3 (b) should appear around the red line but appear in the upper triangular region.
@@ 5 Methods
Figure 4: The framework of STAR.
@@ 5.2 Dynamic Uncertainty Measurement
where $\mu_{b}$ and $\mu_{f}$ denote the prediction uncertainty of the base model and the full model respectively, $\lambda(t)\in[0,1]$ is a monotone decreasing function of AL iteration $t$.
@@ 5.3 Calibration with Hybrid Regularization
[…]
@@ 5.4 Overall Algorithm
5:   querying subset $S_{k}^{U}$ from $D_{k}^{U}$ based on Equation 2
@@
9:   fine-tuning LLM $M_{k}$ to get $M_{k+1}$ on $D_{k+1}^{L}$ based on Equation 6 and Equation 7
@@ 6.2 Settings
Table 2: The ablation performance of different methods, the AUC and RIPL are reported.
@@ 7.1 Main Result
Then, we explore how the models’ performance changes as the training set increases. Figure 5 shows the learning curves for corresponding AL methods on GSM8K, BoolQ, and OpenBookQA datasets, respectively. It is apparent that the Random baseline and the two original active learning approaches perform comparably, suggesting that the active learning methods appear to be ineffective. Notably, the BoolQ dataset exhibits particularly high variability in results when using the PE strategy, which may be attributed to BoolQ’s binary output format of “true” and “false”. The gap between the full model and the base model could easily lead to skewed predicting results in a single iteration.
@@ 7.2 Ablation Study
Since our methods have two main components, which are dynamic uncertain measurement and calibration with hybrid regularization as described in Sec 5. As shown in Table 2, we conduct a detailed ablation study to assess the effect of the two components. Firstly, upon employing the dynamic uncertain measurement, all AUC are enhanced, and the RIPL turns entirely positive. This indicates a significant gap between the full model and the base model in the original strategies, which our dynamic indicator effectively mitigated.
@@ Appendix A Dataset
Dataset
Fine-tuning Data Template
GSM8K
[QUESTION]
Answer the above question. First think step by step and then answer the final number.
[ANSWER]
BoolQ
[QUESTION]
The correct answer is
[ANSWER]
OpenBookQA
[QUESTION]
Answer1: [ANSWER_1]
Answer2: [ANSWER_2]
Answer3: [ANSWER_3]
Answer4: [ANSWER_4]
The correct answer is [ANSWER]

### v2

@@ Abstract
Though Large Language Models (LLMs) have demonstrated the powerful capabilities of few-shot learning through prompting methods, supervised training is still necessary for complex reasoning tasks. Because of their extensive parameters and memory consumption, both Parameter-Efficient Fine-Tuning (PEFT) methods and Memory-Efficient Fine-Tuning methods have been proposed for LLMs. Nevertheless, the issue of large annotated data consumption, the aim of Data-Efficient Fine-Tuning, remains unexplored. One obvious way is to combine the PEFT method with active learning. However, the experimental results show that such a combination is not trivial and yields inferior results. Through probe experiments, such observation might be explained by two main reasons: uncertainty gap and poor model calibration. Therefore, in this paper, we propose a novel approach to effectively integrate uncertainty-based active learning and Low-Rank Adaptation (LoRA). Specifically, for the uncertainty gap, we introduce a dynamic uncertainty measurement that combines the uncertainty of the base model and the uncertainty of the full model during the iteration of active learning. For poor model calibration, we incorporate the regularization method during LoRA training to keep the model from being over-confident, and the Monte-Carlo dropout mechanism is employed to enhance the uncertainty estimation. Experimental results show that the proposed approach outperforms existing baseline models on three complex reasoning tasks.111 Our code and results will be available at https://github.com/callanwu/STAR.
@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Wei et al. (2021); OpenAI (2022); Touvron et al. (2023a, b); Zhao et al. (2023) have demonstrated the powerful capabilities of zero/few-shot learning with prompting techniques, including In-Context Learning Dong et al. (2022) and Chain-of-Thought Wei et al. (2022), where no parameter update is required. However, previous studies Hendrycks et al. (2020); Yuan et al. (2023); Bai et al. (2023); Isik et al. (2024) have shown that further fine-tuning is still crucial for tasks involving complex reasoning such as arithmetic reasoning Roy and Roth (2016); Cobbe et al. (2021) and commonsense reasoning Mihaylov et al. (2018); Clark et al. (2019), etc.
@@
Fine-tuning LLMs requires updating a large number of parameters, which takes a lot of time and consumes considerable memory. Taking LLaMA-7B Touvron et al. (2023a) as an example, fine-tuning it on a dataset of 52k instances takes over 12 hours on 4 A100 80G GPUs Bommasani et al. (2021); Taori et al. (2023). Therefore, Parameter Efficient Fine-Tuning (PEFT) methods  Houlsby et al. (2019); Lester et al. (2021); Li and Liang (2021); Hu et al. (2021); Ding et al. (2023) and Memory Efficient Fine-Tuning (MEFT) methods Liao et al. (2023) have been proposed. In addition to updating a vast number of parameters and consuming substantial memory, a neglected factor in LLMs fine-tuning is the extensive consumption of annotation data. Moreover, due to the inherent complexity of tasks, the human annotation resources required for fine-tuning LLMs are also significant Ouyang et al. (2022).
Therefore, it is important to develop the Data-Efficient Fine-Tuning (DEFT) method for LLMs. A common practice to improve data efficiency is active learning Cohn et al. (1996); Settles (2009), while it has been shown that the PEFT methods can alleviate the reliance on annotated data to some extent Ding et al. (2023). A straightforward idea for DEFT is to combine the PEFT method with active learning. However, such a combination is not trivial. As shown in Figure 1 (a), simply fine-tuning an LLM with LoRA Hu et al. (2021) under an uncertainty-based active learning framework yields consistently inferior performances compared to passive learning (random selection of data in active learning) on the OpenBookQA dataset, while as shown in Figure 1 (b), fine-tuning LLM with full parameter under active learning yields performances better than passive learning.
@@ 2.1 Efficient Fine-tuning Methods
As LLMs continue to expand in size, the computational and financial resources required for fine-tuning these models become increasingly prohibitive. To address this challenge, Efficient Fine-Tuning has emerged as an essential area of research Wan et al. (2023). The methods can be classified into PEFT and MEFT Wan et al. (2023); Liao et al. (2023). PEFT, in particular, aims to adjust a minimal subset of the model’s parameters, thus conserving computational resources while maintaining or enhancing model performance Hu et al. (2023). We classify PEFT methods into four categories: Prompt Tuning Lester et al. (2021); Liu et al. (2023b) only allow an additional $k$ tunable tokens per downstream task to be prepended to the input text. Prefix Tuning Li and Liang (2021); Liu et al. (2022) keeps language model parameters frozen but optimizes a small continuous task-specific vector that pretends to be key-value pairs. Adapter Houlsby et al. (2019); He et al. (2021) is a new module added between layers of a pre-trained network, which is a bottleneck architecture. Low-Rank Adaptation (LoRA) Aghajanyan et al. (2021); Hu et al. (2021) reduces the parameters and enhances computational efficiency by applying low-rank matrices. Moreover, minimizing memory usage in fine-tuning for improving efficiency has also emerged as a critical topic Liao et al. (2023), with several innovative solutions being proposed. Among these, techniques such as QLoRA Dettmers et al. (2023), QA-LoRA Xu et al. (2023), and LoftQ Li et al. (2023) stand out for their ability to significantly reduce memory requirements without compromising model performance. In this paper, we focus on the application of LoRA.
@@ 2.2 Active Learning with LLMs
Active Learning (AL) has been extensively investigated across a multitude of NLP tasks, encompassing machine translation Miura et al. (2016); Zhao et al. (2020), natural language inference Snijders et al. (2023), named entity recognition Shen et al. (2017) and text classification Ein-Dor et al. (2020); Margatina et al. (2022); Schröder et al. (2023). In the era of LLMs, active learning is primarily employed in the selection of prompts and the annotation of data Zhang et al. (2023b); Liu et al. (2023a); Xiao et al. (2023). For instance,  Margatina et al. (2023) explores various active learning strategies for selecting the most relevant examples for in-context learning with LLMs.  Diao et al. (2023) introduces an active prompting method that leverages uncertainty metrics to select questions for annotation. In the domain of integrating PEFT with AL, Jukić and Snajder (2023a) explored PEFT methods with different active learning in Pre-trained language models (PLMs) such as BERT Devlin et al. (2019), demonstrating that the integration of PEFT with active learning can offer substantial performance gains. Different from Jukić and Snajder (2023a), we apply decoder-only generative LLMs as the backbones, to our knowledge, we are the first to integrate LLMs combined PEFT with AL within the realm of reasoning tasks.
@@ 4.1 Probe Experiment Design
We mainly focus on two training variants of LLMs. The first one, denoted as PEFT method, is LLM finetuned with LoRA, which is shown to be problematic under the AL framework. The second one, denoted as Few-shot method, is untuned LLM, which is taken as a control group. To enhance the performance of untuned LLM on downstream tasks, we employ In-Context Learning Dong et al. (2022) prompting by adding demonstrations with the input prompt. LLaMA-2 Touvron et al. (2023b) serves as the backbone LLM. Experiments are conducted on the BoolQ dataset Clark et al. (2019) because its labels only include “true” and “false”, which makes the prediction uncertainty and prediction confidence easy to calculate.
@@ 4.2 Probing with Prediction Confidence
where $p_{true}$ and $p_{false}$ denotes the probabilities of token “true” and “false”, respectively.
Then the prediction confidence of PEFT and Few-shot are calculated and the density plot is drawn to make a comparison between these two methods as shown in Figure 2. To mitigate differences in model accuracy, we only consider the confidence of the model for the wrong predictions on the test set of BoolQ. The intuition is that for examples that the model is less likely to predict right, they should be less confident.
@@ 4.3 Probing with Prediction Entropy
Figure 3: (a) Heatmap of correlation between prediction entropy across different iterations; (b) Scatter plot for prediction entropy between base model (Iter0) and model after first iteration (Iter1); (c) Same as (b), except values are taken from Iter5 and Iter6.
@@
The correlation between examples with top 1000 entropy at the beginning is calculated and the heatmap of the correlation is shown in Figure 3 (a). As we can observe in Figure 3 (a), the correlation between the base model (model without PEFT tuning) and models after AL iteration is close to 0, which indicates a clear gap between the base model and PEFT model. This phenomenon is even clear with the scatter plot in Figure 3, where the dots in Figure 3 (b) should appear around the red line but appear in the upper triangular region. In Figure 3 (c), the correlation coefficients of entropy between the two iterations become relatively normal, which is consistent with Figure 3 (a), suggesting that the gap between iterations has been alleviated.
@@ 5 Methods
Figure 4: The framework of STAR. It primarily consists of five steps: Model Inference, Data Querying, Data Labeling, Dataset Updating, and Model Training.
@@ 5.2 Dynamic Uncertainty Measurement
where $\mu_{b}$ and $\mu_{f}$ denote the prediction uncertainty of the base model and the full model respectively, $\lambda(t)\in[0,1]$ is a monotone decreasing function of AL iteration $t$. Note that, our measurement approach only requires one additional computation of the base model at the beginning, which remains constant throughout and does not significantly increase the FLOPs.
@@ 5.3 Calibration with Hybrid Regularization
$\begin{split}\mu_{f}&=\frac{1}{K}\sum_{k}\mu_{f}^{(k)}\\ \mu_{f}^{(k)}&=\textit{ME}(\text{LLM}(x|\hat{A}_{k},\hat{B}_{k}))\end{split}$ (7)
@@ 5.4 Overall Algorithm
5:   querying subset $S_{k}^{U}$ from $D_{k}^{U}$ based on Equation (2)
@@
9:   fine-tuning LLM $M_{k}$ to get $M_{k+1}$ on $D_{k+1}^{L}$ based on Equation (6) and Equation (7)
@@ 6.2 Settings
Table 2: The ablation performance of different methods, AUC and RIPL are reported.
@@ 7.1 Main Result
Then, we explore how the models’ performance changes as the training set increases. Figure 5 shows the learning curves for corresponding AL methods on GSM8K, BoolQ, and OpenBookQA datasets, respectively. The Random baseline and the two original active learning approaches perform comparably, suggesting that the active learning methods appear to be ineffective. Notably, the BoolQ dataset exhibits particularly high variability in results when using the PE strategy, which may be attributed to BoolQ’s binary output format of “true” and “false”. The gap between the full model and the base model could easily lead to skewed predicting results in a single iteration.
@@ 7.2 Ablation Study
Since our methods have two main components, which are dynamic uncertain measurement and calibration with hybrid regularization as described in Sec 5. We conduct a detailed ablation study to assess the effect of the two components. As shown in Table 2, upon employing the dynamic uncertain measurement, all AUC are improved, and the RIPL turns positive. This indicates a significant gap between the full model and the base model in the original strategies, which our dynamic indicator effectively mitigated.
@@ 8 Conclusion
Acknowledgement
The authors would like to thank the anonymous reviewers for their insightful comments. This work is funded by the National Natural Science Foundation of China (62176053). This work is supported by the Big Data Computing Center of Southeast University.
@@ Appendix A Dataset
Dataset Fine-tuning Data Template
GSM8K [QUESTION] Answer the above question. First, think step by step and then answer the final number. [ANSWER]
BoolQ [QUESTION] The correct answer is [ANSWER]
OpenBookQA [QUESTION] Answer1: [ANSWER_1] Answer2: [ANSWER_2] Answer3: [ANSWER_3] Answer4: [ANSWER_4] The correct answer is [ANSWER]